# CNSocialDepress: A Chinese Social Media Dataset for Depression Risk Detection and Structured Analysis

### v1

@@ Abstract
Depression is a pressing global public health issue, yet publicly available Chinese-language resources for risk detection remain scarce and are mostly limited to binary classification. To address this limitation, we release CNSocialDepress, a benchmark dataset for depression risk detection from Chinese social media posts. The dataset contains 44,178 texts from 233 users, within which psychological experts annotated 10,306 depression-related segments. CNSocialDepress provides binary risk labels together with structured multi-dimensional psychological attributes, enabling interpretable and fine-grained analysis of depressive signals. Experimental results demonstrate its utility across a wide range of NLP tasks, including structured psychological profiling and fine-tuning of large language models for depression detection. Comprehensive evaluations highlight the dataset’s effectiveness and practical value for depression risk identification and psychological analysis, thereby providing insights to mental health applications tailored for Chinese-speaking populations.
@@
Jinyuan XU1,*, Tian LAN2,*, Xintao YU3, Xue HE3,4, Hezhi ZHANG5, Ying WANG6, Pierre Magistry1, Mathieu Valette1, Lei LI7,8†thanks: * These authors contributed equally to this work.thanks: † Corresponding Author. 1 Ertim Inalco, 2 Milkuya Studio, 3 Sorbonne Université, 4 IRD Lab 5 Faculty of Psychology, Peking University 6 Faculty of Psychology and Cognitive Science, Beijing Normal University 7 University of Washington, 8 VitaSight
1 Introduction
@@
Depressive disorders are among the most common mental health conditions worldwide, characterized by persistent low mood or loss of interest in daily activities. According to a 2023 report [^1] https://www.who.int/news-room/fact-sheets/detail/depression by the World Health Organization (WHO), approximately 280 million people globally suffer from depression. In China, a 2024 survey by the Chinese Center for Disease Control and Prevention Wang et al. (2024a) estimates that around 95 million individuals are affected by depression, and out of the 280,000 suicides reported annually, 40% are linked to depressive disorders. Moreover, research Wang et al. (2024a) shows that depression strongly correlates with both suicidal and non-suicidal self-injurious deaths.
@@
Motivated by the urgent need for the detection of depression, researchers increasingly apply machine learning (ML) and natural language processing (NLP) methods for automated analysis of depression risk Squires et al. (2023); Hasib et al. (2023); Aleem et al. (2022). While early efforts have shown promising results, they are heavily constrained by the limitations of existing datasets. Traditional depression detection studies often rely on clinical dataBittar et al. (2019); Fernandes et al. (2018), or on transcripts from medical or psychological interviews resources Shen et al. (2022); Li et al. (2022) that are are expensive to collect, limited in size and diversity, and may not reflect the informal, emotionally nuanced expressions typical of real-world online environments.
@@
To overcome these limitations, recent research has shifted toward leveraging user-generated content on social media platforms [^2] https://github.com/bucuram/depression-datasets-nlp Harrigian et al. (2021), which provides rich linguistic signals and greater availability. Such data are also more easily anonymized, alleviating some privacy concerns. Beyond these methodological advantages, social media itself has become an emotional outlet for many individuals, especially in East Asian contexts Yang and Li (2009); Zhou et al. (2023), where personal emotions are more freely expressed online rather than in face to face settings. Social media serves as a critical outlet for emotional expression, potentially offering richer and more authentic linguistic data for depression analysis.
@@
However, most existing datasets are in English and focus solely on classification with binary or multi-class labels, lacking structured psychological insight or professional validation. Cai et al. (2023); Wang et al. (2020); Li et al. (2020). Moreover, with the emergence of generative models applied to mental health Hu et al. (2024); Xu et al. (2024); Lai et al. (2023); Yang et al. (2024), many scholars in psychology and computational linguistics now emphasize the need for datasets combining risk labels with structured analyses, such as user profiles or survey-based explanations. Such datasets align more closely with real-world settings, where multi-faceted insights are helpful for clinical intervention.
@@
Some recent efforts have introduced summarization style datasets for depression, leveraging information extraction and automated summarization techniques Sotudeh et al. (2022, 2021). However, these datasets usually rely on model generated content validated by automatic metrics (e.g., ROUGE Lin (2004)) and typically lack professional mental health annotation, raising concerns about domain-specific reliability and accuracy.
@@
To address these challenges, we introduce (CNSocialDepress, the first publicly available Chinese-language dataset for depression risk detection that integrates binary risk labels with structured psychological analyses. All annotations are drafted and validated by certified mental health experts, ensuring both domain relevance and annotation quality. CNSD is designed to support multiple task paradigms, including binary classification, generation of structured analyses, summarization, and fine-tuning of large language models (LLMs) for psychological reasoning.
@@
  • We construct a high-quality Chinese dataset that bridges the gap between classification and interpretable analysis in depression detection, and supports a wide range of tasks including binary classification, generation of structured psychological analyses, summarization, and fine-tuning of large language models.
  • We design a robust annotation and generation pipeline involving professional experts and structured templates;
  • We conduct extensive experiments demonstrating that CNSD enables effective training and evaluation of LLM-based methods across diverse depression-related tasks.
@@
  • We propose and validate a robust pipeline for generating structured psychological analyses in the context of depression risk (Section 4).
2 Related Work
2.1 Existing Datasets for Depression Detection
@@
Current depression detection datasets primarily come from English-language social media platforms like Twitter, Reddit, Facebook, and Instagram Shen et al. (2017); Parapar et al. (2022); Zhang et al. (2021); Raihan et al. (2024). Annotation methods include self-disclosure identification Yates et al. (2017); Bathina et al. (2021); Islam et al. (2018), manual coding by clinical raters Almouzini et al. (2019); Yazdavar et al. (2020); Alhamed et al. (2024), and symptom mapping Seabrook et al. (2018); Aldarwish and Ahmad (2017); Zhang et al. (2022) based on DSM American Psychiatric Association et al. (2000) and PHQ Kroenke et al. (2001). While most datasets use binary classification, others incorporate severity scales or symptom-specific tagging Zhang et al. (2022); Mowery et al. (2017). Multilingual resources for depression detection exist for Spanish Romero et al. (2024), Arabic Maghraby and Ali (2022), Russian Stankevich et al. (2020), Portuguese Santos et al. (2024), Japanese Yuka Niimi (2021), and Thai Hämäläinen et al. (2021). For Chinese, Sina Weibo is the primary source Li et al. (2020); Shen et al. (2018); Li et al. (2023); Guo et al. (2023); Yang et al. (2021), with datasets like WU3D Wang et al. (2020) and SWDD Cai et al. (2023) commonly used.
2.2 Methodological Advancements
Early methods in depression detection used statistical approaches like TF-IDF and feature engineering Yang et al. (2020) along with traditional machine learning classifiers Cortes (1995); Li (2024); Breiman (2001); McCallum et al. (1998); Dreiseitl and Ohno-Machado (2002); Li et al. (2024); Shi et al. (2024). With the rise of neural networks Schmidhuber (2015), deep learning models such as Word2Vec Mikolov et al. (2013), LSTM Sak et al. (2014), CNN Kim (2014), Transformer Vaswani (2017), and BERT Devlin et al. (2019) improved the detection of depression. Recently, large language models (LLMs) have been applied to depression detection tasks due to their ability to generate labels and provide interpretable analyses Lan et al. (2024); Hu et al. (2024); Xin Yan (2023); Lai et al. (2023). These models enhance explainability in mental health analysis Wang et al. (2024c); Yang et al. (2024); Xu et al. (2024); Yang et al. (2023); Hu et al. (2024); Xu et al. (2024).
2.3 Summary and Gaps
@@
Despite existing efforts, current datasets rarely include structured psychological analysis, especially for non-English social media content. Furthermore, available resources often lack professional mental-health annotation, limiting their practical utility. This study addresses these limitations by presenting the first expert-validated Chinese-language dataset featuring detailed structured annotations, thereby providing richer resources for depression-risk detection and analysis.
3 Construction of Dataset
Figure 1: Dataset Construction Process: During the data annotation process, we used a subset of the original SWDD dataset for labeling, which included 116 Positive (Depressive) users and 117 Negative users. Psychologists used rating scales based on the DSM-5 and the statistical results of the dataset’s text to formulate an initial labeling guideline. Then, during the psychologists’ labeling process, random sampling was continuously performed for cross validation, and the labeling standards were continuously updated based on the annotation outcomes. Ultimately, the labeled gold-standard data was obtained, with each user containing a six-dimensional structural analyses summary.
Our raw data are derived from the SWDD dataset Cai et al. (2023), a user-level dataset collected from Sina Weibo, the world’s largest Chinese-language social media platform. This dataset comprises two categories of users (depressed and non-depressed), each user has between dozens and hundreds of Weibo posts. The dataset, which is labeled as either depressed or non-depressed, also contains user-level depression feature annotations provided by experts in accordance with the DSM-5333https://www.psychiatry.org/psychiatrists/practice/dsm/educational-resources/dsm-5-fact-sheets American Psychiatric Association et al. (2000) reference standards.
@@ 3 Construction of Dataset
From this SWDD dataset, we have 116 self-reported depressed (positive) users and 117 non-depressed (negative) users to form the candidate samples for our dataset. Each selected user has at least 60 posts, and the total token count for each user’s posts is no fewer than 3,000 tokens [^4] Qwen2.5’s tokenizer.
@@
To guide the annotation process, a group of psychology experts formulated a preliminary annotation standard based on the DSM-5 diagnostic criteria, the PHQ-9 scale Kroenke et al. (2001), and prior research on linguistic features indicative of depression in online texts Mothe et al. (2022).
@@
Criteria involving objective factual statements (e.g., clinical symptoms, medical records, self-reports) with stronger diagnostic specificity are designated as Primary Criteria, while criteria based on subjective linguistic or emotional expressions (e.g., negative phrasing, emotional intensity) with weaker specificity are designated as lower-weight Secondary Criteria, and the hierarchical classification follows clinical diagnostic principles that prioritize objective symptoms over subjective perceptions. This standard comprises six depression dimensions:
@@
  • Dimension 1: Depressive Psychological State (primary criterion). At its core, this category reflects a loss of self-worth, overwhelming guilt, and suicidal ideation. Representative semantic segments include Inferiority, Sorry, Death, Self-harm, Wrist cutting, Suicide, etc.
@@
  • Dimension 2: Depression-Related Medical Expressions (primary criterion). Focusing on expressions linked to medication usage and psychological diagnoses, this category covers terms such as Take medication, Venlafaxine, Side effects, Depression, Anxiety disorder, Hospital, Doctor, Disease, etc.
@@
  • Dimension 3: Depression-Related Clinical Symptoms (primary criterion). Addressing physiological and somatic alterations, including changes in weight or appetite, disturbed sleep patterns, fatigue, and various forms of physical pain, this dimension includes semantic segments such as Appetite, Insomnia, Sleeping pills, Nightmare, Tired, Headache, Stomachache, among others.
  • Dimension 4: Negative Emotions (secondary criterion). Encompassing a broad array of adverse emotional states, this part presents examples such as Torment, Abyss, Hell, Sad, Grieve, Anxious, Depressed, Lonely, Despair, etc.
@@
  • Dimension 5: Potential External Causes of Depression (secondary criterion). Highlighting possible triggers like marital or intimate relationship issues, family conflicts, significant social events, and everyday stressors, this segment comprises semantic units including Divorce, Parents, College entrance examination, School, Teacher, Poor, Drop out, etc.
@@
  • Dimension 6: Depression-Related Patterns of Language Use (secondary criterion). Emphasizing particular linguistic patterns, such as negations, interrogatives, and derogatory expressions. This category includes illustrative semantic segments like I don’t know, I don’t like me, I really don’t want to, I can’t do it, why, and what should I do?, among others.
@@
Figure 2: Example entry from the CNSD-Gold dataset
The annotation work was carried out by four senior researchers, each with substantial research experience in psychology, specializing in depression scales. Using an internally developed annotation platform, they identified specific depression dimensions present in each post, along with the corresponding semantic segments. Throughout this process, the annotators periodically performed random cross-validation to ensure reliability and consistency. Because this annotation task relies on qualitative rather than quantitative metrics, the annotation schema was updated as needed, informed by real-time feedback from the annotated texts and discussions among experts. During the annotation process, the expert annotators did not refer to the binary labels from the original dataset, consequently, some of the depression risk labels may differ from those in the original data set.
@@
This rigorous process produced the CNSD Gold dataset (100 positive and 100 negative users) and the CNSD Test dataset (16 positive and 17 negative users). The dataset construction process can be referenced in Figure 1.
@@
Each entry as shown in Figure 2 comprises two parts: a binary depression label and a detailed, dimension-specific analysis. For each dimension, the analysis includes an overall assessment and a list of post numbers with brief justifications for their selection.
@@
The overall statistics are presented in Table 1, with a total of 10,306 dimension-level annotations summarized in Table 2.
@@
Number of Users 116 117
Number of Texts 20,360 23,818
Number of Tokens 1,024,978 1,115,951
@@
Table 2: Dimension Annotation Statistics for Negative and Positive Users: This table shows the statistics of depression repeated segments marked by psychological experts across different dimensions. A text may contain multiple depression repeated segments.
4 Automated Dataset Generation Pipeline
Due to the substantial human effort and professional psychological expertise required for dataset annotation, user-level depression risk datasets based on social media remain scarce in the Chinese domain. To address this gap, we developed an automated dataset generation pipeline. This pipeline was used to create the CNSD Silver dataset by processing the SWDD binary classification dataset, leveraging insights from the manually annotated CNSD Gold dataset (described in Section 3). For the construction of CNSD Silver, we randomly selected 100 positive (depressive risk) users and 100 negative (non-depressive risk) users from the SWDD dataset. Each user contributed between several dozen and several hundred posts, totaling more than 3,000 tokens per user.
The pipeline consists of two main modules:
4.1 Module I: Dimension-Wise Automatic Labeling
This module uses a smaller model (Qwen2.5-14B Bai et al. (2023)) to automatically annotate texts across six depression-related dimensions ($\mathcal{D}=\{D_{1},D_{2},\dots,D_{6}\}$). The model is first fine-tuned on carefully curated text-level data derived from the positive users in CNSD Gold, enabling it to assign each individual text to one or more of these six dimensions.
Data Preparation for Fine-tuning Module I Model:

### v2

@@ Abstract
Depression is a pressing global public health issue, yet publicly available Chinese-language resources for depression risk detection remain scarce and largely focus on binary classification. To address this limitation, we release CNSocialDepress, a benchmark dataset for depression risk detection on Chinese social media. The dataset contains 44,178 posts from 233 users; psychological experts annotated 10,306 depression-related segments. CNSocialDepress provides binary risk labels along with structured, multidimensional psychological attributes, enabling interpretable and fine-grained analyses of depressive signals. Experimental results demonstrate the dataset’s utility across a range of NLP tasks, including structured psychological profiling and fine-tuning large language models for depression detection. Comprehensive evaluations highlight the dataset’s effectiveness and practical value for depression risk identification and psychological analysis, thereby providing insights for mental health applications tailored to Chinese-speaking populations.
Keywords: Depression Detection, Chinese Social Media, Benchmark Dataset, Mental Health
[…]
@@
Jinyuan XU1*, Tian LAN2*, Xintao YU3, Xue HE3,4, Hezhi ZHANG5, Ying WANG6 Pierre Magistry1, Mathieu Valette1, Lei LI7†
1Ertim Inalco, 2Milkuya Studio, 3Sorbonne Université, 4IRD Lab,
5Faculty of Psychology, Peking University,
6Faculty of Psychology and Cognitive Science, Beijing Normal University,
7Beijing Institute of Technology
Abstract content
11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding Author.
1.  Introduction
@@
Depressive disorders are among the most common mental health conditions worldwide and are characterized by persistent low mood or loss of interest in daily activities. According to a 2023 report [^1] https://www.who.int/news-room/fact-sheets/detail/depression by the World Health Organization (WHO), approximately 280 million people worldwide live with depression. In China, a 2024 survey by the Chinese Center for Disease Control and Prevention Wang et al. (2024a) estimates that around 95 million individuals are affected by depression. Of the approximately 280,000 suicides reported annually in China, 40% are linked to depressive disorders. Moreover, research Wang et al. (2024a) shows that depression is strongly associated with both suicidal behavior and non-suicidal self-injury.
@@
Motivated by the urgent need to detect depression, researchers increasingly apply machine learning (ML) and natural language processing (NLP) methods to automatically assess depression risk Squires et al. (2023); Hasib et al. (2023); Aleem et al. (2022); Liu et al. (2024); Shi et al. (2024); Jia and Li (2024). While early efforts have shown promising results, they remain constrained by the limitations of existing datasets. Traditional depression detection studies often rely on clinical data Bittar et al. (2019); Fernandes et al. (2018) or on transcripts from medical or psychological interviews Shen et al. (2022); Li et al. (2022b), which are expensive to collect, limited in size and diversity, and may not reflect the informal, emotionally nuanced expressions typical of real-world online environments.
@@
To overcome these limitations, recent research has shifted toward leveraging user-generated content on social media platforms Bucur et al. (2025); Harrigian et al. (2021); Li et al. (2025d); Jiang et al. (2024); Cai et al. (2025a), which provides rich linguistic signals and is more readily available. Such data are also more easily anonymized, thereby alleviating some privacy concerns. Beyond these methodological advantages, social media has become an important channel for emotional expression, especially in East Asian contexts Zhou et al. (2023); Yang and Li (2009); Zhou et al. (2024); Yang et al. (2025), where personal emotions may be expressed more freely online than in face-to-face settings. As a result, social media can provide richer and more authentic linguistic signals for depression analysis.
@@
However, most existing datasets focus solely on classification with binary or multi-class labels and lack structured psychological insights or professional validation. Moreover, generative models are increasingly used in mental health applications Hu et al. (2024); Xu et al. (2024); Yang et al. (2024b); Lai et al. (2023); He and Qu (2025a); Gu et al. (2025); Cai et al. (2025b); Xu et al. (2025). Researchers in psychology and computational linguistics emphasize the need for datasets that combine risk labels with structured analyses, such as user profiles or survey-based explanations. Such datasets better match real-world settings. Multifaceted insights can support clinical decision-making and downstream intervention.
@@
Some recent efforts have introduced summarization-style datasets for depression. They leverage information extraction and automated summarization techniques Sotudeh et al. (2022, 2021); He and Qu (2025b); Shi et al. . However, these datasets often rely on model-generated content. The content is typically validated using automatic metrics such as ROUGE Lin (2004). They also lack annotation by mental health professionals, which raises concerns about domain-specific reliability and accuracy.
@@
To address these challenges, we introduce CNSocialDepress (CNSD), a publicly available Chinese-language dataset for depression risk detection that can be applied to the development of early-stage depression detection tools. It pairs binary risk labels with structured psychological analyses. All annotations are drafted and validated by certified mental health professionals, ensuring domain relevance and annotation quality. CNSD supports multiple task paradigms. These include binary classification, structured analysis generation, summarization, and fine-tuning large language models (LLMs) for psychological reasoning.
@@
  • We release CNSD, a high-quality Chinese dataset for depression risk detection that combines binary labels with structured, interpretable psychological analyses.
  • We propose an expert-in-the-loop annotation protocol with structured templates and quality control.
  • We benchmark CNSD across diverse task settings, including classification, structured analysis generation, summarization, and LLM fine-tuning for psychological reasoning.
@@
  • We present and validate a pipeline for generating structured psychological analyses for depression risk (Section 4).
2.  Related Work
2.1.  Existing Datasets for Depression Detection
@@
Current datasets for depression detection mainly come from English-language social media platforms such as Twitter, Reddit, Facebook, and Instagram Shen et al. (2017); Parapar et al. (2022); Zhang et al. (2021); Raihan et al. (2024); Li et al. (2025c); Jiang et al. (2025). Common annotation strategies include identifying self-disclosure Yates et al. (2017); Bathina et al. (2021); Islam et al. (2018), manual coding by clinical raters Almouzini et al. (2019); Yazdavar et al. (2020); Alhamed et al. (2024), and symptom mapping Seabrook et al. (2018); Aldarwish and Ahmad (2017); Zhang et al. (2022); Li et al. (2025a). Symptom definitions are often based on DSM-5 criteria Association (2013) and questionnaire instruments such as the PHQ-9 Kroenke et al. (2001).
Most datasets use binary labels. Others incorporate severity scales or symptom-specific tags Zhang et al. (2022); Mowery et al. (2017); Guan et al. (2025). Multilingual resources also exist for depression detection. They cover languages such as Spanish Romero et al. (2024), Arabic Maghraby and Ali (2022), Russian Stankevich et al. (2020), Portuguese Santos et al. (2024), Japanese Yuka Niimi (2021), and Thai Hämäläinen et al. (2021). For Chinese, Sina Weibo is the most common data source Li et al. (2020); Shen et al. (2018); Li et al. (2023); Guo et al. (2023); Yang et al. (2021); Li et al. (2025b). Datasets such as WU3D Wang et al. (2020a) and SWDD Cai et al. (2023) are widely used.
$2.$  Methodological Advancements
Early work relied on statistical representations such as TF–IDF and hand-crafted features Yang et al. (2020); Li et al. (2022a). These features were typically paired with traditional machine learning classifiers Cortes (1995); Breiman (2001); McCallum et al. (1998); Dreiseitl and Ohno-Machado (2002); He et al. (2026). With the rise of neural approaches Schmidhuber (2015), representation learning became central to depression detection. Embedding methods such as Word2Vec Mikolov et al. (2013) and neural architectures including LSTMs Sak et al. (2014), CNNs Kim (2014), Transformers Vaswani (2017), and BERT Devlin et al. (2019) improved performance.
Recently, large language models (LLMs) have been explored for depression detection Lan et al. (2024); Hu et al. (2024); Xin Yan (2023); Lai et al. (2023); Li et al. (2026); He and Qu (2025a); Shi et al. (2025). They can produce predictions and generate natural-language rationales or structured analyses, which supports interpretability. LLM-based approaches have also been used to improve explainability in mental health analysis Wang et al. (2024c); Yang et al. (2024b); Xu et al. (2024); Yang et al. (2023); Hu et al. (2024); Gu et al. (2025).
2.3.  Summary and Gaps
@@
Despite these efforts, few datasets include structured psychological analyses, especially for non-English social media. Furthermore, available resources often lack annotation by mental health professionals, which limits practical utility. This study addresses these limitations by presenting an expert-validated Chinese-language dataset with detailed structured annotations. The dataset provides richer resources for depression risk detection and analysis.
3.  Construction of Dataset
Figure 1: Dataset construction process. We sample 116 positive (depressed) users and 117 negative users from SWDD for expert annotation. Psychologists draft an initial guideline based on DSM-5 and corpus statistics, and iteratively refine it through random cross-validation. The final gold-standard data include a six-dimensional structured analysis summary for each user.
Our raw data come from the SWDD dataset Cai et al. (2023), a user-level corpus collected from Sina Weibo, one of the largest Chinese-language social media platforms. The dataset includes two user groups: depressed and non-depressed. Each user has dozens to hundreds of Weibo posts. In addition to binary user labels, SWDD provides expert-annotated user-level depression features based on DSM-5 criteria [^2] DSM-5 Fact Sheets. Association (2013).
@@ 3.  Construction of Dataset
From SWDD, we select 116 self-reported depressed (positive) users who disclosed a clinical diagnosis of depression in their posts, and 117 non-depressed (negative) users as candidates. Each selected user has at least 60 posts. The total length of a user’s posts is at least 3,000 tokens, computed using the Qwen2.5 tokenizer. [^3] Qwen2.5 tokenizer.
@@
To guide annotation, a team of psychology experts drafted an initial guideline based on DSM-5 criteria, the PHQ-9 Kroenke et al. (2001), and prior work on linguistic markers of depression in online text Mothe et al. (2022).
@@
We define two levels of criteria. Primary criteria cover objective statements with higher diagnostic specificity, such as clinical symptoms, medical records, and explicit self-reports. Secondary criteria capture subjective linguistic or emotional expressions with lower specificity, such as negative phrasing and emotional intensity. This hierarchy follows clinical principles that prioritize objective symptoms over subjective perceptions. The guidelines cover six dimensions:
@@
  • Dimension 1: Depressive psychological state (primary). This dimension reflects loss of self-worth, overwhelming guilt, and suicidal ideation. Representative spans include inferiority, apology, death, self-harm, and suicide.
@@
  • Dimension 2: Medical expressions related to depression (primary). This dimension covers mentions of medication use and clinical diagnoses, such as taking medication, venlafaxine, side effects, depression, anxiety disorder, hospital, and doctor.
@@
  • Dimension 3: Clinical symptoms related to depression (primary). This dimension captures physiological or somatic changes, including appetite or weight changes, sleep disturbance, fatigue, and physical pain. Representative spans include appetite, insomnia, sleeping pills, nightmares, tiredness, headache, and stomachache.
  • Dimension 4: Negative emotions (secondary). This dimension covers adverse emotional states such as sadness, grief, anxiety, loneliness, and despair.
@@
  • Dimension 5: Potential external causes of depression (secondary). This dimension captures possible triggers such as intimate relationship issues, family conflict, major social events, and everyday stressors. Representative spans include divorce, parents, school, teacher, and dropping out.
@@
  • Dimension 6: Language Use Patterns Related to Depression (secondary). This dimension captures linguistic patterns such as negation, questions, and derogatory expressions. Representative spans include I don’t know, I don’t like myself, I don’t want to, I can’t do it, why, and what should I do ?.
@@
Figure 2: A user-level annotated entry from CNSD-Gold (English translations).
Each user has multiple posts. We treat each post as the primary unit for review. Annotation is performed on semantic spans within each post. Each span is assigned to one of the six dimensions. A post may contain multiple spans, and different spans may belong to different dimensions.
Annotation was conducted by four senior researchers with expertise in psychology and depression scales. Annotators were blinded to the original binary labels in SWDD to ensure independent judgment. They periodically cross-checked overlapping samples to maintain consistency. Because the task relied more on qualitative judgment than on fixed quantitative criteria, the annotation guideline was not set in stone from the outset. It was revised in rounds as experts reviewed newly labeled data and resolved disagreements through discussion, and earlier annotations were corrected when the updated standards called for it. Given this evolving process, conventional inter-annotator agreement scores are not directly applicable. Since our annotation follows a different scheme from SWDD and all labeling was performed independently of the original labels, some user-level labels in our dataset differ from those in the original dataset.
@@
This process yields the CNSD Gold set (100 positive and 100 negative users) and the CNSD Test set (16 positive and 17 negative users). Figure 1 summarizes the construction process.
@@
Each entry (Figure 2) contains a binary label and a dimension-specific analysis. For each dimension, the analysis provides an overall assessment. It also lists post indices with brief justifications.
@@
The overall statistics are presented in Table 1, and the 10,306 dimension-level annotations are summarized in Table 2.
@@
NO. of Users 116 117
NO. of Texts 20,360 23,818
NO. of Tokens 1,024,978 1,115,951
@@
Table 2: Dimension-level annotation statistics for negative and positive users. This table reports the number of expert-annotated depression-related spans in each dimension. A single text may contain multiple spans.
4.  Automated Dataset Generation Pipeline
Expert annotation requires substantial human effort and professional psychological expertise. As a result, user-level depression-risk datasets for Chinese social media remain scarce. To narrow this gap, we develop an automated dataset generation pipeline. We use it to construct CNSD Silver from the SWDD binary classification dataset. The pipeline is informed by the manually annotated CNSD Gold dataset (Section 3).
To build CNSD Silver, we randomly sample 100 positive (depression-risk) users and 100 negative (non-risk) users from SWDD. Each user contributes dozens to hundreds of posts. We keep users with at least 3,000 tokens in total.
The pipeline consists of two modules.
4.1.  Module I: Dimension-Wise Automatic Labeling
Module I uses a mid-sized model (Qwen2.5-14B Bai et al. (2023)) to automatically assign posts to the six depression-related dimensions ($\mathcal{D}=\{D_{1},D_{2},\dots,D_{6}\}$). We fine-tune the model on text-level training data curated from positive users in CNSD Gold. After fine-tuning, the model can assign each post to one or more dimensions.
Data preparation for fine-tuning Module I.